two cards are drawn from a well shuffled pack of cards . the probability that one is a heart card and the other is king

Grade:12

1 Answers

Arun
6 years ago
Dear Nithin
 

Let X be the event that cards are in a Heart which is not king and other is the king of heart.
Let Y be the event that one is any heart card and other is a non-heart king.
Hence, required probability:
=P(A)+P(B)

=\inline {\color{Black}[(12C_{1}\times 1C_{1})/52C_{2}]+[(13C_{1}\times 3C_{1})/52C_{2}]}

 on solving

 

= 1/26

 

Regards

Arun (askIITians forum expert)